OpenStudy (lina17):

solve for x: -4|x+5|=-16

OpenStudy (4everaddicted2anime):

First you need to divide both sides by -4

OpenStudy (lina17):

|x+5|=4

OpenStudy (4everaddicted2anime):

Now you need to find an x value that'll add with 5 to get either a positive or negative 4

OpenStudy (lina17):

so do i add

OpenStudy (4everaddicted2anime):

Subtract 5 from each side

OpenStudy (leozap1):

x=−1,−9

OpenStudy (lina17):

im confused

alones (alones):

do you want to see my steps again :)

OpenStudy (lina17):

yes

OpenStudy (leozap1):

I'll show you step by step, just a sec.

OpenStudy (leozap1):

Divide each term in −4|x+5|=−16 by 4 to eliminate the fractions. −|x+5|=−4

alones (alones):

okay so tep #1: Divide both sides by -4. −4(|x+5|)−4=−16−4 |x+5|=4 Step 2: Solve Absolute Value. |x+5|=4 We know eitherx+5=4orx+5=−4 x+5=4(Possibility 1) x+5−5=4−5(Subtract 5 from both sides) x=−1 x+5=−4(Possibility 2) x+5−5=−4−5(Subtract 5 from both sides) x=−9 Answer: x=−1 or x=−9 do you understand??

OpenStudy (leozap1):

Divide each term in −|x+5|=−4 by −1 to eliminate the fractions. |x+5|=4

OpenStudy (lina17):

thank you so much @AloneS

OpenStudy (leozap1):

Remove the absolute value term. This creates a ± on the right-hand side of the equation because |x|=±x. x+5=±(4)

OpenStudy (leozap1):

Set up the positive portion of the ± solution. x+5=4

OpenStudy (leozap1):

Move all terms not containing x to the right-hand side of the equation. x=−1

OpenStudy (leozap1):

Set up the negative portion of the ± solution. x+5=−(4)

OpenStudy (leozap1):

Simplify the right-hand side. x+5=−4

alones (alones):

You welcome

OpenStudy (leozap1):

Move all terms not containing x to the right-hand side of the equation. x=−9

OpenStudy (leozap1):

The solution to the equation includes both the positive and negative portions of the solution. x=−1,−9